Mavis ate 3/10
of her birthday cake. She decided to give the brother 2/3
of what is left of the cake.
1
Expert's answer
2020-09-08T16:13:44-0400

Mavis ate "\frac{3}{10}" of the birthday cake


i.e "1- \frac{3}{10}" "= \frac{7}{10}"


"\frac{2}{3}" what was left of the cake she decided to give her brother, i.e.


"\frac{7}{10}×\frac{2}{3}" "= \frac{7}{15}" was eaten by her brother,

"\frac{7}{10}" was left and brother ate "\frac{7}{15}" .


So "\frac{7}{10}-\frac{7}{15}=\frac{7}{30}" was left.